# A class of $d$-dimensional regular black holes: Shadows, Thermodynamics and Gravitational collapse

### v1

@@ 7 Conclusion
This paper has focused on the class of $4$-dimensional regular BHs with a de Sitter core, which was recently shown to be generated by the collapse of a polytropic star [11]. This class of BHs contains the well-known Bardeen and Hayward regular BHs. We extended this class to the $d$-dimensional case and considered nonlinear electrodynamics as the source of Einstein’s equations. The presence of magnetic charge in the source results in a regular center, whereas its absence results in an ST BH. Our class reduces to $d$-dimensional Bardeen and Hayward BHs when special values are chosen for the polytropic index.
We have analyzed the horizons, energy conditions, extremal conditions, photon spheres, and shadows of the aforementioned class of BHs. We have demonstrated that the SEC is violated at certain radii and, depending on the polytropic index value, the DEC may also be violated for specific dimensions and radii. Our findings show that, as the number of dimensions increases, the differences in the radii of horizons, photon spheres, and shadows decrease. Photon spheres and shadows persist only up to a limit of magnetic charge which we call $q_{\text{deg}}$. Beyond this limit, the photon spheres and shadow vanish, resulting in a horizonless compact object. Additionally, the extremality condition is independent of the polytropic index. Therefore, it is the same for both Bardeen and Hayward BHs, as well as for other regular BHs belonging to the mentioned class of BHs in $4$ dimensions.
We have studied the thermodynamics of our proposed class of $d$-dimensional regular BHs using the standard definitions of surface gravity, temperature, entropy, and heat capacity. Unlike the ST BH, zero temperature occurs at a minimal horizon radius, i.e, the extremal horizon, $r_{\text{ext}}$. Horizon radii below this value are considered unphysical because they correspond to negative temperature. As the horizon radius increases from $r_{\text{ext}}$, the temperature rises from zero to a maximum, after which it decreases for larger horizon radii.
We have also examined the heat capacity and thermodynamic stability of the aforementioned BH model. Unlike an ST BH, which is always unstable, our proposed $d$-dimensional regular BHs can be thermodynamically stable for certain horizon radii. They also exhibit possible phase transitions for every dimension and polytropic index.
Furthermore, we have generalized the OSD collapse to $d$ dimensions and replaced the Schwarzschild exterior with a $d$-dimensional regular BH solution. We have analyzed the evolution of the stellar surface, as well as the apparent and event horizons. As the number of dimensions increases, we have found that the formation of a BH takes a longer proper time. Increasing the polytropic index slightly increases the proper time for collapse as well.

### v2

@@ 7 Conclusion
In this work, we have conducted a comprehensive investigation into a general class of $d$-dimensional regular BHs characterized by a de Sitter core, sourced by nonlinear electrodynamics. This framework serves as a higher-dimensional generalization of the well-known 4-dimensional Bardeen and Hayward BHs, which are recovered for specific values of the polytropic index $n$ [11]. By extending the analysis beyond the standard four-dimensional spacetime, we have elucidated the intricate interplay between the spacetime dimensionality, the magnetic charge parameter, and the polytropic index in shaping the geometric, optical, thermodynamic, and dynamic properties of regular BHs.
Our analysis of the metric structure reveals that the presence of magnetic charge ensures a regular center with finite curvature, thereby resolving the central singularity inherent in classical solutions. In the absence of charge, the solution reduces to the singular ST BH. We derived exact conditions for the existence of horizons and demonstrated that while the extremality condition for the charge is independent of the polytropic index, the extremal radius depends on it. A key geometric finding is that as the number of dimensions increases, the distinctions between the radii of the event horizons, photon spheres, and shadows diminish, suggesting that higher-dimensional regular BHs converge behaviorally towards their singular counterparts in terms of geometric scales. Furthermore, we analyzed the ECs satisfied by the source matter. While the WEC and NEC are satisfied everywhere, the SEC is violated at certain radii, and the DEC may also be violated depending on the specific values of $n$ and $d$, indicating the presence of exotic matter regions necessary to sustain the regular core.
Regarding the optical properties, we identified the existence of both stable and unstable photon spheres. For BH configurations, an unstable outer photon sphere exists, whereas horizonless compact objects (formed when $q>q_{\text{ext}}$) can possess both stable and unstable photon spheres. We established that photon spheres and shadows persist only up to a critical magnetic charge limit, denoted as $q_{\text{deg}}$. Beyond this threshold, the object transitions into a horizonless compact object devoid of a shadow. Our numerical results indicate that as the dimension $d$, the magnetic charge $q$, or the polytropic index $n$ increases, the shadow’s size decreases. Crucially, we observed that as the number of dimensions increases, the differences between the radius of the photon sphere and the radius of the shadow become negligible, implying that detailed studies of photon spheres in very high dimensions may be less critical for observational distinctions.
Thermodynamically, the proposed regular BHs exhibit rich behavior distinct from the ST BH. Unlike the ST BH, which is always thermodynamically unstable, our model possesses regions of local stability characterized by a positive heat capacity. We observed that the temperature vanishes at the extremal horizon radius, in contrast to the ST BH, for which the temperature diverges as the horizon radius approaches zero. Therefore, the extremal horizon radius must be imposed as the lower limit in the calculations of entropy and heat capacity, as it has not been considered in previous studies. The heat capacity exhibits singularities signaling possible phase transitions, and the entropy deviates from the standard area law. Notably, this deviation from the area law becomes more significant as the dimension increases, while increasing the polytropic index reduces the deviation. These features suggest that regular BHs offer a more consistent thermodynamic description in the context of quantum gravity corrections, particularly in higher dimensions where the entropy-area relation is modified.
Finally, we generalized the OSD collapse scenario to this $d$-dimensional regular background. By matching an interior FLRW metric to the exterior regular solution, we tracked the evolution of the stellar surface, apparent horizon, and event horizon analytically. Our results indicate that the formation of the BH takes a longer proper time as the number of dimensions and the polytropic index increase. We also derived a critical lower bound for the initial stellar radius ($R_{0,\text{min}}$) required for physical BH formation. Through a perturbative analysis for small magnetic charges, we demonstrated that small values of the magnetic charge decrease this minimum initial radius, thereby facilitating the collapse scenario compared to the neutral case. Specifically, in 4 dimensions, we found explicit corrections for the Hayward and Bardeen cases, showing that the reduction in $R_{0,\text{min}}$ is sensitive to the polytropic index.
Collectively, these results provide new insights into the viability of regular BHs as non-singular endpoints of gravitational collapse in higher dimensions. The dependence of stability, shadow size, and collapse dynamics on the polytropic index and spacetime dimension offers potential observational signatures that could distinguish these regular solutions from classical BHs in future astrophysical surveys. Future work could explore the observational signatures of these $d$-dimensional objects in the context of gravitational wave astronomy and higher-dimensional holographic models.
@@ A.3 Horizons in $9$ dimensions
Appendix B A hypergeometric identity
Here we detail the simplification leading to Eq. (123). The starting point is the differentiation of $\tau_{B}$ with respect to $R_{0}$ using Eq. (6.1)
$\frac{d\tau_{B}}{dR_{0}}=\frac{d}{dR_{0}}\left[\frac{2m^{\frac{3-d}{2}}}{d-1}\left(R_{0}^{\frac{d-1}{2}}\,{}_{2}F_{1}\!\left[a,a;1-a;-cR_{0}^{-b}\right]-R_{B}^{\frac{d-1}{2}}\,{}_{2}F_{1}\left[a,a;1-a;-cR_{B}^{-b}\right]\right)\right],$ (B.1)
where
$a\equiv-\frac{(d-1)n}{3(d-2)},\quad b\equiv\frac{3(d-2)}{2n},\quad c\equiv m^{\frac{d-2}{2}(\frac{3}{n}-2)}q^{d-2}.$ (B.2)
Applying the hypergeometric derivative identity
$\frac{d}{dz}\,{}_{2}F_{1}[a,a;1-a;z]=\frac{a^{2}}{1-a}{}_{2}F_{1}[a+1,a+1;2-a;z],\quad z\equiv-cR_{0}^{-b}.$ (B.3)
together with $dz/dR_{0}=cbR_{0}^{-b-1}$, we obtain
$\frac{d\tau_{B}}{dR_{0}}$ $=\frac{2m^{\frac{3-d}{2}}}{d-1}\left[\frac{d-1}{2}R_{0}^{\frac{d-3}{2}}\,{}_{2}F_{1}[a,a;1-a;z]+zbR_{0}^{\frac{d-3}{2}}\frac{a^{2}}{1-a}\,{}_{2}F_{1}[a+1,a+1;2-a;z]\right],$ (B.4)
$=\frac{2m^{\frac{3-d}{2}}}{d-1}R_{0}^{\frac{d-3}{2}}\left[\frac{d-1}{2}\left({}_{2}F_{1}[a,a;1-a;z]+\frac{2zb}{d-1}\frac{a^{2}}{1-a}\,{}_{2}F_{1}[a+1,a+1;2-a;z]\right)\right],$ (B.5)
$=\frac{R_{0}^{\frac{d-3}{2}}}{m^{\frac{d-3}{2}}}\left[{}_{2}F_{1}[a,a;1-a;z]+\frac{az}{1-a}\,{}_{2}F_{1}[a+1,a+1;2-a;z]\right].$ (B.6)
where in the last step we used $\frac{2b}{d-1}\frac{a^{2}}{1-a}=-\frac{a}{1-a}$. To further simplify the expression, we use the series representations. For the first term we have
${}_{2}F_{1}[a,a;1-a;z]=\sum_{k=0}^{\infty}\frac{(a)_{k}(a)_{k}}{(1-a)_{k}}\frac{z^{k}}{k!},$ (B.7)
where $(a)_{k}$ is Pochhammer symbol. The second term in (B.6) can be written as
$z\,{}_{2}F_{1}[a+1,a+1;2-a;z]=z\sum_{k=0}^{\infty}\frac{(a+1)_{k}(a+1)_{k}}{(2-a)_{k}}\frac{z^{k}}{k!}.$ (B.8)
Shifting the summation index $k\to k-1$ in the second term and using Pochhammer identities, $(a+1)_{k-1}=\frac{(a)_{k}}{a}$ and $(2-a)_{k-1}=\frac{(1-a)_{k}}{1-a}$ yields
$\frac{az}{1-a}\,{}_{2}F_{1}[a+1,a+1;2-a;z]=\sum_{k=1}^{\infty}\frac{(a)_{k}(a+1)_{k-1}}{(1-a)_{k}}\frac{z^{k}}{(k-1)!}.$ (B.9)
Thus,
${}_{2}F_{1}[a,a;1-a;z]+\frac{a}{1-a}\,z\,{}_{2}F_{1}[a+1,a+1;2-a;z]$ $=\sum_{k=0}^{\infty}\frac{(a)_{k}^{2}}{(1-a)_{k}}\frac{z^{k}}{k!}+\sum_{k=1}^{\infty}\frac{(a)_{k}(a+1)_{k-1}}{(1-a)_{k}}\frac{kz^{k}}{k!},$ (B.10)
$=1+\sum_{k=1}^{\infty}\frac{(a)_{k}}{(1-a)_{k}}\frac{z^{k}}{k!}\left[(a)_{k}+k(a+1)_{k-1}\right].$ (B.11)
Using the identity $(a)_{k}+k(a+1)_{k-1}=(a+1)_{k}$ we obtain the compact hypergeometric form
${}_{2}F_{1}[a,a;1-a;z]+\frac{a}{1-a}\,z\,{}_{2}F_{1}[a+1,a+1;2-a;z]$ $=\sum_{k=0}^{\infty}\frac{(a)_{k}(a+1)_{k}}{(1-a)_{k}}\frac{z^{k}}{k!}$ (B.12)
$={}_{2}F_{1}[a,a+1;1-a;z]$ (B.13)
Substituting back $a$, $z$ (and subsequently $c$ and $b$) yields (123).